Use the COR (Coefficient Of Restitution) to answer the following:

A baseball is launched at a bat with a speed of 95 mph. The baseball rebounded with a speed of 58 mph. What is the COR of this baseball?

A basketball should have a minimum COR of .82. If a "short" referee drops the ball from a height of 1600 mm, how high should the ball bounce after hitting the floor. Keep your answer in millimeters.
Answer #1

Coefficient of restitution = 58/95 = 0.61

--------------------------------------------------

velocity v of the ball at ground, if it is dropped from 1600 mm is given by

v = [ 2 g H ]1/2 = [ 2 9.8 1.6 ]1/2 = 5.6 m/s

if Coefficient of restitution is 0.82, then rebound velocity = 5.6 0.82 m/s 4.6 m/s

height h raised by the ball is given by ,

h = [ v2 / (2g) ]1/2 = [ ( 4.6 4.6 ) / ( 29.8) ]1/2 = 1.039 m = 1039 mm